# A Robust Large-Period Discrete Time Crystal and its Signature in a Digital Quantum Computer

### v4

@@ 1 Introduction
The concept of non-ergodicity [1] in quantum phenomena is ubiquitous and important in quantum many-body physics [2]. It underlies a variety of exotic physical phenomena such as eigenstate thermalization hypothesis [3, 4, 5], many-body localization [6, 7], quantum scars [8, 9], quantum chaos [10], and time crystals [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. In particular, the time crystals are further categorized into two classes, i.e., the continuous time crystals (CTCs) and discrete time crystals (DTCs). The CTCs more closely follow the original proposal of Ref. [11] in which continuous time-translational symmetry is spontaneously broken. The CTCs have recently been implemented in various mesoscopic-scale experimental platforms [16, 20, 18, 21, 22, 23]. The second class, i.e., the DTCs [24, 25], are a type of non-ergodic phases of matter  [^1] Strictly speaking, a phase of matter is characterized by macroscopic behaviors that survive in the thermodynamic limit. that gained prominence in the recent years [26, 27]. The DTCs emerge in periodically driven systems and are characterized by the presence of an order parameter evolving at a period that is robustly locked at an integer multiple of the driving period, persisting indefinitely in the thermodynamic limit [14, 15, 26, 27].
@@
Experimentally realizing DTCs whose order parameter exhibits a much larger period than the corresponding driving period is highly desirable, as it paves the way for observing passive quantum error correction [28], as well as novel dynamical physics such as Anderson localization and Mott insulator transitions in the time domain [29, 30, 31, 32, 33, 34, 35]. Unfortunately, despite a considerable number of theoretical proposals for such large period DTCs [36, 37, 38, 39, 40, 41, 42], a majority of existing experiments were only able to realize period-doubling [43, 44, 45, 46, 47, 48] and period-tripling [49] DTCs. Indeed, as these experiments utilize (pseudo) spin-1/2 particles or spin-1 particles respectively, their underlying $\mathbb{Z}_{2}$ or $\mathbb{Z}_{3}$ symmetry only natively supports period-doubling or period tripling feature respectively. By contrast, the theoretical proposals of Refs. [36, 37, 38, 39] utilize higher-spin bosonic particles, which are thus incompatible with these experiments. Nevertheless, it is worth pointing out that experimental demonstrations of large period DTCs and even their quasicrystal generalizations have recently been made in systems of higher-spin bosonic particles, such as Refs. [39, 16, 50]. However, to the best of our knowledge, experimental realizations of large period DTCs in systems containing strictly two-level particles are still lacking. On the other hand, two-level particles (qubits) form the most fundamental building blocks of a typical quantum processor. Meanwhile, realizing quantum phases of matter in a quantum processor has been an active research area in recent years, as it paves the way towards establishing quantum advantage that eventually leads to the development of a functional quantum computer.
@@
In this work, we develop an interacting spin-1/2 system that supports period-quadrupling DTCs (which we shall refer to as $4T$-DTCs) observable even at moderate system sizes. The time-evolution with matrix product states (tMPS) [52, 53] method enables us to numerically investigate, in a controlled manner, the effect of disorder which is necessarily present in the actual quantum devices. Indeed, existing noisy intermediate-scale quantum (NISQ) devices currently possess various kinds of noise, ranging from the relatively poor gate fidelity, the deep circuit depth, to the thermal environment noise rising from the execution of the quantum circuit [54, 55, 56]. In this case, the disorder analysis we carry out in the numerical studies simulates the imperfect gate fidelity of NISQ devices. Remarkably, we found that the signatures of $4T$-DTCs are not only robust against various types of disorders/gate errors, but can even be amplified in some cases.
Finally, in the efforts towards simulating quantum many-body systems on superconducting quantum processors [57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69], we implement our proposed system in the IBM Q quantum processor ibmq_cairo. Despite the various aforementioned noise occurring in our NISQ-era device, a robust period-quadrupling order parameter are clearly observed even at the accessible moderate number of qubits. Our successful observation and analysis of period-quadrupling DTCs in a NISQ-era device are expected to further bridge the gap between fundamental studies of DTCs and their technological applications
2 Results
2.1 $4T$-Discrete Time Crystal
@@
2.2 Evidence of robust $4T$-DTC via tMPS
@@
Figure 2: Numerical evidence of robust $4T$-DTC for $N=16$ sites using tMPS. (a) Magnetization $\langle S_{z}\rangle$ as a function of time at $MT=0.98\pi$, which is slightly perturbed away from the “ideal" solvable limit value $MT=\pi$. The 4T-DTC phase (red triangles) corresponds to $hT=0.9\pi$ (which is also perturbed from $\pi$) and $JT=0.16\pi$. The thermal phase (filled orange squares) corresponds to $hT=0.52\pi$ and $JT=0.1\pi$. The $JT=0$ phase (empty blue squares) corresponds to $hT=0.9\pi$. (b) The associated stroboscopic power spectrum $\langle\tilde{S}_{z}\rangle$, which shows distinct frequency peaks at $\Omega=\pm\pi/2T$ only for the 4T-DTC phase.
@@
where $\mathcal{N}_{\rm tot}$ is the total stroboscopic steps evolved, and $t=kT$ ($k=1,2,\cdots,\mathcal{N}_{\rm tot}$) is the stroboscopic time at step $k$. Our results are summarized in Fig. 2 for two different sets of parameter values that correspond to the $4T$-DTC phase and the thermal phase respectively. Specifically, as the parameters $h$ and $M$ are chosen close to but not equal to the solvable limit values, at a finite value of the intersite interaction $J$, the oscillatory feature of $\langle S_{z}\rangle$ is clearly observed (triangle markers in Fig. 2(a)). That this oscillatory feature indeed corresponds to period-quadrupling behavior is further demonstrated by a sharp peak at the subharmonic frequency components $\Omega=\pi/2,3\pi/2$ in the power spectrum of Fig. 2(b). Moreover, we find that such a $4T$-periodicity is observed over a window of parameter values and not only at a specific set of parameter values, which suggests that the system indeed supports a $4T$-DTC phase. If a parameter $h$ or $M$ deviates significantly from its corresponding ideal value, $\langle S_{z}\rangle$ quickly decays to zero, and the system is in the thermal phase (empty square markers in Fig. 2(a)). If the intersite interaction $J$ is absent, $\langle S_{z}\rangle$ may seem to exhibit finite amplitudes according to the filled square markers in Fig. 2(a). However, the power spectrum plot of Fig. 2(b) reveals that, similarly to the case of thermal phase, there are no significant peaks at any frequency, thereby demonstrating the absence of period-quadrupling behavior. Any remnant of oscillatory behavior in the cases of $JT=0$ and thermal phase as observed in Fig. 2(a) is attributed to the fact that our system Hamiltonian has the tendency to flip some “spin" states over the course of one period, as again illustrated in the schematics of Fig. 1(b).
@@
In Fig. 3(a), we obtain the phase diagram of the noise-free system by plotting the subharmonic $\Omega T=\pi/2$ peak ($\langle\tilde{S}_{z}\rangle\left(\Omega\right)$ at $\Omega=\pi/2T$) in the power spectrum against the two system parameters $h$ and $J$ As the value $\Omega=\pi/2T$ represents the fundamental Fourier component of a periodic variable with frequency $\pi/2T$, it is expected that a profile exhibiting a sharp peak at $\Omega=\pi/2T$ is close to being $4T$-periodic. We can then associate a finite (zero) $\langle\tilde{S}_{z}\rangle\left(\Omega\right)$ at $\Omega=\pi/2T$ with the $4T$-DTC (thermal) phase, of which the amplitude quantifies the $4T$-oscillation signature. It is observed that the $4T$-DTC phase spans over a considerable window of $h$ values – symmetrically about $hT=\pi$–at moderate values of $J$. At $J=0$, the period-quadrupling feature is observed only at $hT=\pi$, further confirming the role of the intersite interaction in stabilizing the DTC phase. On the other hand, at very large values of $J$, the $4T$-DTC behavior is absent altogether, which could be attributed to the presence of quantum chaos [74].
@@
2.3 Finite size scaling analysis of the spectral properties of the many-body Floquet eigenstates
In Sec. 2.1, we have analytically elucidated the origin of our 4T-DTC model in terms of the spectral structure of the many-body Floquet eigenstates, i.e., they are long-range correlated and exhibit $\pi/2$-spectral gap. These features are exact and clearly seen at ideal parameter values whereby all many-body Floquet eigenstates can be analytically shown to take the form of macroscopic $4$-component cat states. Away from these ideal values, we argued that such features due to a mechanism reminiscent of quantum error correction. To further verify this argument and therefore the validity of $4T$-DTC, we shall numerically perform a thorough finite size scaling analysis of the aforementioned spectral properties in the following.
@@ 2.3 Finite size scaling analysis of the spectral properties of the many-body Floquet eigenstates
The case $\chi_{zz}=1$ signifies perfect long-range correlation with respect to $ZZ$ Pauli operator, whereas the other extreme case of $\chi_{zz}=0$ implies its absence. Our calculation (as shown in Fig. 4(a) (red triangle) at generic parameters deep in the DTC region) reveals that $\chi_{zz}$ takes an intermediate value between 0 and $1$, and it monotonically increases with the system size, thereby demonstrating the presence of long-range correlation in the many-body Floquet eigenstates with respect to $ZZ$ Pauli operator in the thermodynamic limit.
Figure 4: Finite size scaling analysis as a validity check of $4T$-DTC. (a) The Ising-ZZZZ spin-glass (SG) order $\chi_{zz}$ and the $\pi/2$-shifted spectral function $s_{\pi/2}$ as a function of the total system size; (b) The full power spectrum for different sizes from $N=4$ to $12$. All results are obtained by exact diagonalization of the Floquet unitary operator for the time evolution of the total system up to $t=400T$. In all panels, other parameters used are: $JT=0.16\pi,MT=1.0\pi$, and $hT=0.99\pi$.
@@
and $\hat{O}$ is an appropriately chosen operator that is expected to map a many-body Floquet eigenstate to another that is separated in quasienergy by $\pi/(2T)$. In our calculation, we have chosen $\hat{O}=\sigma_{a,0}^{z}+i\sigma_{b,0}^{z}$ as it yields the exact desired mapping at ideal parameter values. Indeed, at ideal parameter values ($J=0$ and $hT=\pi$), $s_{\pi/2}$ can be analytically computed to yield precisely $1$, which indicates that all many-body Floquet eigenstates exhibit perfect $\pi/2$ spectral gap. At more general parameter values, our numerical calculation of Fig. 4(a) (blue circles) reveals that $s_{\pi/2}$ takes a value between 0 and $1$, and it increases monotonically with the system size. This suggests that the many-body Floquet eigenstates indeed have the tendency to exhibit $\pi/2$-shifted spectral gap in the thermodynamic limit, and our operator choice $\hat{O}$ is relatively effective at capturing this spectral gap.
@@
Finally, we have also performed a thorough analysis of the full power spectrum, calculated over a considerably long evolution time of up to $400T$, for various system sizes in Fig. 4(b). As expected, the subharmonic peaks at $1/4$ and $3/4$ the driving frequency increase monotonically with the system size, thereby highlighting the lifetime improvement of the period-quadrupling behavior with the system size.
@@
To conclude, our numerics have confirmed that our system represents a genuine $4T$-DTC based on its three defining characteristics that persist at general parameter values: (i) the many-body Floquet eigenstates are long-range correlated ($\chi_{zz}$ is finite and increases monotonically with the system size), (ii) the corresponding Floquet quasienergies exhibit $\pi/2$-shifted spectral gap ($s_{\pi/2}$ is finite and increases monotonically with the system size), and (iii) the period-quadrupling behavior lasts longer with increase in the system size (the subharmonic peaks at $1/4$ and $3/4$ the driving frequency increase monotonically with the system size).
2.4 Realization of robust $4T$-DTC on a quantum processor

### v5

@@ I Introduction
The concept of non-ergodicity [1] in quantum phenomena is ubiquitous and important in quantum many-body physics [2]. It underlies a variety of exotic physical phenomena such as eigenstate thermalization hypothesis [3, 4, 5], many-body localization [6, 7], quantum scars [8, 9], quantum chaos [10], and time crystals [11, 12, 13, 14, 15].In particular, the time crystals are further categorized into two classes, i.e., the continuous time crystals (CTCs) [16, 17, 18, 19, 20] and discrete time crystals (DTCs). The CTCs more closely follow the original proposal of Ref. [11] in which continuous time-translational symmetry is spontaneously broken. The CTCs (including time quasi-crystals [21]) have recently been implemented in various mesoscopic-scale experimental platforms [19, 20, 16, 17, 18]. The second class, i.e., the DTCs, originally proposed in [22, 23], are a type of non-ergodic phases of matter  [^1] Strictly speaking, a phase of matter is characterized by macroscopic behaviors that survive in the thermodynamic limit. that gained prominence in the recent years [25, 26, 27, 28, 29]. The DTCs emerge in periodically driven systems and are characterized by the presence of an order parameter evolving at a period that is robustly locked at an integer multiple of the driving period, persisting indefinitely in the thermodynamic limit [14, 15, 26, 27].
@@
Experimentally realizing DTCs whose order parameter exhibits a much larger period than the corresponding driving period is highly desirable, as it paves the way for observing passive quantum error correction [30], as well as novel dynamical physics such as Anderson localization [31, 32] and Mott insulator transitions in the time domain [33, 34, 35, 36, 37]. Unfortunately, despite a considerable number of theoretical proposals for such large period DTCs in ultacold atomic systems [38, 39, 40] and in spin-based systems [41, 42, 43, 44], a majority of existing experiments were only able to realize period-doubling DTCs, such as Refs. [45, 46] which utilize superconducting quantum processors or Refs. [47, 48, 49, 50, 51] in other quantum simulation platforms, period-tripling [52] DTCs, and most recently, discrete time quasicrystals [53]. Indeed, as these experiments utilize (pseudo) spin-1/2 particles or spin-1 particles respectively, their underlying $\mathbb{Z}_{2}$ or $\mathbb{Z}_{3}$ symmetry only natively supports period-doubling or period tripling feature respectively. By contrast, the theoretical proposals of Refs. [38, 41, 39, 40] utilize higher-spin bosonic particles, which are thus incompatible with these experiments. Nevertheless, it is worth pointing out that experimental demonstrations of large period DTCs and even their quasicrystal generalizations have recently been made in systems of higher-spin bosonic particles, such as Refs. [40, 21, 54]. However, to the best of our knowledge, experimental realizations of large period DTCs in systems containing strictly two-level particles are still lacking. On the other hand, two-level particles (qubits) form the most fundamental building blocks of a typical quantum processor. Meanwhile, realizing quantum phases of matter in a quantum processor has been an active research area in recent years, as it paves the way towards establishing quantum advantage that eventually leads to the development of a functional quantum computer.
@@
In this work, we develop an interacting spin-1/2 system that supports period-quadrupling DTCs (which we shall refer to as $4T$-DTCs) observable even at moderate system sizes. Intuitively, this is made by possible by carefully designing the periodic driving scheme such that the system exhibits an emergent $\mathbb{Z}_{4}$ symmetry, which is required for the formation of $4T$-DTCs, even though its building blocks (two-level particles) only support $\mathbb{Z}_{2}$ symmetry. The time-evolution with matrix product states (tMPS) [56, 57] method then enables us to numerically investigate, in a controlled manner, the effect of disorder which is necessarily present in the actual quantum devices. Indeed, existing noisy intermediate-scale quantum (NISQ) devices currently possess various kinds of noise, ranging from the relatively poor gate fidelity, the deep circuit depth, to the thermal environment noise rising from the execution of the quantum circuit [58, 59, 60]. In this case, the disorder analysis we carry out in the numerical studies simulates the imperfect gate fidelity of NISQ devices. Remarkably, we found that the signatures of $4T$-DTCs are not only robust against various types of disorders, gate errors and perturbations in the initial state, but can even be amplified in some cases.
Finally, in the efforts towards simulating many-body condensed matter systems [61, 62, 63], including capturing the non-equilibrium quantum phenomena [64, 65, 45, 46, 66], strongly-correlated quantum systems [67, 68, 69] and topological quantum physics [70, 71] on superconducting quantum processors, particularly for realizing DTCs beyond the period-doubling class previously achieved in Refs. [45, 46], we implement our proposed system in the IBM Q quantum processor ibmq_cairo. Despite the various aforementioned noise occurring in our NISQ-era device, a robust period-quadrupling order parameter are clearly observed even at the accessible moderate number of qubits. Thus, our successful observation and analysis of period-quadrupling DTCs in a NISQ-era device are expected to further bridge the gap between fundamental studies of DTCs and their technological applications. In particular, achieving a large-size higher-period DTC beyond the period-doubling type in a system of qubits is not just a quantitative change since generating the necessary $\mathbb{Z}_{n}$ symmetry from the native $\mathbb{Z}_{2}$ symmetry of qubits is a highly nontrivial task. Moreover, it could lead to quantum technological applications, such as for realizing a robust quantum memory or a passive quantum error-correcting device. The generalizability of our nontrivial DTC construction based on two-level particles also provides a fertile avenue for future theoretical and experimental studies to uncover rich phenomenology in the intersection of driven quantum matter and many-body quantum simulations on a hardware.
II Results
II.1 $4T$-Discrete Time Crystal
@@
II.2 Evidence of robust $4T$-DTC via tMPS
@@
Figure 2: Numerical evidence of robust $4T$-DTC for $N=16$ sites using tMPS. (a) Magnetization $\langle S_{z}\rangle$ as a function of time at $MT=0.98\pi$, which is slightly perturbed away from the “ideal” solvable limit value $MT=\pi$. The 4T-DTC phase (red triangles) corresponds to $hT=0.9\pi$ (which is also perturbed from $\pi$) and $JT=0.16\pi$. The thermal phase (filled orange squares) corresponds to $hT=0.52\pi$ and $JT=0.1\pi$. The $JT=0$ phase (empty blue squares) corresponds to $hT=0.9\pi$. (b) The associated stroboscopic power spectrum $\langle\tilde{S}_{z}\rangle$, which shows distinct frequency peaks at $\Omega=\pm\pi/2T$ only for the 4T-DTC phase.
@@
where $\mathcal{N}_{\rm tot}$ is the total stroboscopic steps evolved, and $t=kT$ ($k=1,2,\cdots,\mathcal{N}_{\rm tot}$) is the stroboscopic time at step $k$. Our results are summarized in Fig. 2 for two different sets of parameter values that correspond to the $4T$-DTC phase and the thermal phase respectively. Specifically, as the parameters $h$ and $M$ are chosen close to but not equal to the solvable limit values, at a finite value of the intersite interaction $J$, the oscillatory feature of $\langle S_{z}\rangle$ is clearly observed (triangle markers in Fig. 2(a)). That this oscillatory feature indeed corresponds to period-quadrupling behavior is further demonstrated by a sharp peak at the subharmonic frequency components $\Omega=\pi/2,3\pi/2$ in the power spectrum of Fig. 2(b). Moreover, we find that such a $4T$-periodicity is observed over a window of parameter values and not only at a specific set of parameter values, which suggests that the system indeed supports a $4T$-DTC phase. If a parameter $h$ or $M$ deviates significantly from its corresponding ideal value, $\langle S_{z}\rangle$ quickly decays to zero, and the system is in the thermal phase (empty square markers in Fig. 2(a)). If the intersite interaction $J$ is absent, $\langle S_{z}\rangle$ may seem to exhibit finite amplitudes according to the filled square markers in Fig. 2(a). However, the power spectrum plot of Fig. 2(b) reveals that, similarly to the case of thermal phase, there are no significant peaks at any frequency, thereby demonstrating the absence of period-quadrupling behavior. Any remnant of oscillatory behavior in the cases of $JT=0$ and thermal phase as observed in Fig. 2(a) is attributed to the fact that our system Hamiltonian has the tendency to flip some “spin” states over the course of one period, as again illustrated in the schematics of Fig. 1(b).
@@
In Fig. 3(a), we obtain the phase diagram of the noise-free system by plotting the subharmonic $\Omega T=\pi/2$ peak ($\langle\tilde{S}_{z}\rangle\left(\Omega\right)$ at $\Omega=\pi/2T$) in the power spectrum against the two system parameters $h$ and $J$. As the value $\Omega=\pi/2T$ represents the fundamental Fourier component of a periodic variable with frequency $\pi/2T$, it is expected that a profile exhibiting a sharp peak at $\Omega=\pi/2T$ is close to being $4T$-periodic. We can then associate a finite (zero) $\langle\tilde{S}_{z}\rangle\left(\Omega\right)$ at $\Omega=\pi/2T$ with the $4T$-DTC (thermal) phase, of which the amplitude quantifies the $4T$-oscillation signature. It is observed that the $4T$-DTC phase spans over a considerable window of $h$ values – symmetrically about $hT=\pi$–at moderate values of $J$. At $J=0$, the period-quadrupling feature is observed only at $hT=\pi$, further confirming the role of the intersite interaction in stabilizing the DTC phase. On the other hand, at very large values of $J$, the $4T$-DTC behavior is absent altogether, which could be attributed to the presence of quantum chaos [76].
@@
II.3 Finite size scaling analysis of the spectral properties of the many-body Floquet eigenstates
In Sec. II.1, we have analytically elucidated the origin of our 4T-DTC model in terms of the spectral structure of the many-body Floquet eigenstates, i.e., they are long-range correlated and exhibit $\pi/2$-spectral gap. These features are exact and clearly seen at ideal parameter values whereby all many-body Floquet eigenstates can be analytically shown to take the form of macroscopic $4$-component cat states. Away from these ideal values, we argued that such features due to a mechanism reminiscent of quantum error correction. To further verify this argument and therefore the validity of $4T$-DTC, we shall numerically perform a thorough finite size scaling analysis of the aforementioned spectral properties in the following.
@@ II.3 Finite size scaling analysis of the spectral properties of the many-body Floquet eigenstates
The case $\chi_{zz}=1$ signifies perfect long-range correlation with respect to $ZZ$ Pauli operator, whereas the other extreme case of $\chi_{zz}=0$ implies its absence. Our calculation, as shown in Fig. 4(a) (red triangles) at generic parameters deep in the DTC region, reveals that $\chi_{zz}$ takes an intermediate value between 0 and 11 for $MT=1.0\pi$ and $MT=0.95\pi$. In both cases, $\chi_{zz}$ monotonically increases with the system size, thereby demonstrating the presence of long-range correlation in the many-body Floquet eigenstates with respect to $ZZ$ Pauli operator in the thermodynamic limit. As the same signature is observed for at least two considerably different parameter values, it represents a genuine physical effect and is not a result of some fine-tuning.
Figure 4: Finite size scaling analysis as a validity check of $4T$-DTC. (a) The Ising-ZZZZ SG order $\chi_{zz}$ and the $\pi/2$-shifted spectral function $s_{\pi/2}$ as a function of the total system size for $M=1.0\pi$ and $MT=0.95\pi$. (b) The full power spectrum for different sizes from $N=4$ to $12$ for $MT=0.95\pi$. (c) Comparison of the averaged peak values of the power spectrum for both $MT=1.0\pi$ and $MT=0.95\pi$ as a function of system sizes. All results are obtained by exact diagonalization of the Floquet unitary operator for the time evolution of the total system up to $t=400T$. In all panels, other parameters used are: $JT=0.16\pi$, and $hT=0.99\pi$.
@@
and $\hat{O}$ is an appropriately chosen operator that is expected to map a many-body Floquet eigenstate to another that is separated in quasienergy by $\pi/(2T)$. In our calculation, we have chosen $\hat{O}=\sigma_{a,0}^{z}+i\sigma_{b,0}^{z}$ as it yields the exact desired mapping at ideal parameter values. Indeed, at ideal parameter values ($J=0$ and $hT=\pi$), $s_{\pi/2}$ can be analytically computed to yield precisely $1$, which indicates that all many-body Floquet eigenstates exhibit perfect $\pi/2$ spectral gap. At more general parameter values, our numerical calculation of Fig. 4(a) (blue circles) reveals that $s_{\pi/2}$ takes a value between 0 and $1$, and it increases monotonically with the system size. Similarly, as qualitatively the same signature is observed for two different parameter values $MT=\pi$ and $MT=0.95\pi$, the observed effect is not a result of fine-tuning. Instead, it genuinely demonstrates that the many-body Floquet eigenstates indeed have the tendency to exhibit $\pi/2$-shifted spectral gap in the thermodynamic limit, and our operator choice $\hat{O}$ is relatively effective at capturing this spectral gap.
@@
In addition, we have also performed a thorough analysis of the full power spectrum, calculated over a considerably long evolution time of up to $400T$, for various system sizes in Fig. 4(b) for $MT=0.95\pi$ that is deviated from the perfect spin flipping parameter $MT=1.0\pi$. As expected, the subharmonic peaks at $1/4$ and $3/4$ the driving frequency increase monotonically with the system size, thereby highlighting the lifetime improvement of the period-quadrupling behavior with the system size. While not shown in the figure, qualitatively the same profiles are obtained for $MT=1.0\pi$. To clearly demonstrate this, we have also compared the subharmonic peaks value ($\langle\tilde{S}_{z}\rangle_{\text{max}}$), defined as the average of peak values at $1/4$ and $3/4$, for both $MT=1.0\pi$ and $MT=0.95\pi$ in Fig. 4(c). In both cases, $\langle\tilde{S}_{z}\rangle_{\text{max}}$ clearly increases with the system size, thereby rigorously establishing the lifetime improvement relation with system sizes.
Figure 5: Demonstration of the robustness against different initial state choices. (a,b) the full power spectrum for different impurity initial state parameters $\theta_{y}$ with different system sizes in (a) $N=6$ and (b) $N=12$. (c) The $\pi/2$-shifted spectral function $s_{\pi/2}$ and (d) the Ising-$ZZ$ SG order as a function of system sizes for different impurity initial state parameters $\theta_{y}$.
Finally, we have also examined the robustness of the $4T$-DTCs signatures against variations in the initial state. This is achieved by applying an appropriate single qubit rotation with respect to Pauli $Y$ to each qubit before inputting all qubits to the circuit simulating our system. The resulting operator can be written as $R_{y}\left(\theta_{y}\right)=\prod_{i}\exp\left[-i\Theta_{i}^{y}\frac{\sigma_{i}^{y}}{2}\right]$, where the product is over all qubits involved and $\Theta_{i}^{y}$ is randomly drawn from a uniform distribution $\left[-\theta_{y},\theta_{y}\right]$. In Fig. 5(a,b), we repeated the calculation of power spectrum for two different system sizes and several different angular deviation $\theta_{y}$ values. In all cases, the subharmonic peaks at $1/4$ and $3/4$ are still significantly observed. Moreover, the larger system size $N=12$ yields larger peak values for all $\theta_{y}$ values under consideration, thereby demonstrating the robustness of the $4T$-oscillation against variations in the initial state, as expected from a genuine DTC.
Strictly speaking, the Ising-$ZZ$ SG parameter $\chi_{zz}$ and the $\pi/2$-shifted spectral function $s_{\pi/2}$ capture the spectral features of the system’s Floquet operator and thus do not directly depend on the initial state. However, the effect of varying the initial state could be indirectly imprinted on these quantities by applying the rotating operator (via $R_{y}^{(-1)}(\theta_{y})(\cdots)R_{y}(\theta_{y})$) on the relevant observable, i.e., $(\cdots)\equiv\hat{\mathcal{O}}$ for $s_{\pi/2}$ and $(\cdots)\equiv\sum_{\begin{subarray}{c}i,i^{\prime}\\ \gamma,\gamma^{\prime}\end{subarray}}\sigma_{i,\gamma}^{z}\sigma^{z}_{i^{\prime},\gamma^{\prime}}$ for $\chi_{zz}$. In Fig. 5(c,d), several $\theta_{y}$-modified $\chi_{zz}$ and $s_{\pi/2}$ parameters are plotted with the system size. Remarkably, even when largely deviated from the all spin-up initial state, e.g., with $\theta_{y}=0.2\pi$, non-trivial values for both $s_{\pi/2}$ and $\chi_{zz}$ that improve with increasing system sizes are still clearly observed in Fig. 5(c) and (d) respectively. All these results provide clear evidence that the $4T$-DTCs signatures are robust against variations from the all spin-up initial state.
@@
To conclude, our thorough numerics have confirmed that our system represents a genuine $4T$-DTC based on its three defining characteristics that persist at general parameter values and various initial states: (i) the many-body Floquet eigenstates are long-range correlated ($\chi_{zz}$ is finite and increases monotonically with the system size), (ii) the corresponding Floquet quasienergies exhibit $\pi/2$-shifted spectral gap ($s_{\pi/2}$ is finite and increases monotonically with the system size), (iii) the period-quadrupling behavior lasts longer with increase in the system size (the subharmonic peaks at $1/4$ and $3/4$ the driving frequency increase monotonically with the system size).
II.4 Realization of robust $4T$-DTC on a quantum processor